I am three times as old as my son. Five years later i shall be two and a half times as old as my son.What is my age?
Answer by stanbon(75887)

I am three times as old as my son. Five years later i shall be two and a half times as old as my son.What is my age?
---
Equations:
m = 3s
(m+5) = (5/2)(s+5)
---
Substitute for "m" and solve for "s":
3s+5 = (5/2)s + (25/2)
(1/2)s = 15/2
s = 15 (son's age now)
----
m = 3s = 45 (father's age now)
==================================
cheers,
Stan H.
================